OpenStudy (anonymous):

Which of the following numbers is irrational? negative 7 over 4, 2 over 3, square root of 3, square root of 4 a fraction with numerator negative 7 and denominator 4 a fraction with numerator 2 and denominator 3 square root of 3 square root of 4

OpenStudy (anonymous):

help

OpenStudy (usukidoll):

\[\LARGE -\frac{7}{4}, \frac{2}{3}, \sqrt{3},\sqrt{4}\] like this?

OpenStudy (anonymous):

yes @UsukiDoll

OpenStudy (usukidoll):

ok the definition of a rational number is any number that can be expressed as the quotient or fraction p/q of two integers, p and q, with the denominator q not equal to zero. Since q may be equal to 1, every integer is a rational number.

OpenStudy (usukidoll):

irrational numbers can't be expressed as fractions. having square roots is a dead giveaway of an irrational number

OpenStudy (usukidoll):

an irrational number is any real number that cannot be expressed as a ratio of integers. Irrational numbers cannot be represented as terminating or repeating decimals.

OpenStudy (usukidoll):

\[\LARGE e, \pi, \sqrt{2}\] are examples of irrational numbers.

OpenStudy (usukidoll):

so from this list \[\LARGE -\frac{7}{4}, \frac{2}{3}, \sqrt{3},\sqrt{4} \] which of them are irrational?

OpenStudy (anonymous):

3

OpenStudy (usukidoll):

yes the square root of 3 is irrational

OpenStudy (usukidoll):

the square root of 4 can be simplified further into 2 and it's rational because it can be expressed as a fraction which is \[\LARGE \frac{2}{1}\]

OpenStudy (usukidoll):

\[\LARGE \frac{2}{1} \rightarrow 2 \]

OpenStudy (usukidoll):

everything else is rational \[\LARGE \sqrt{3} \] is irrational.

OpenStudy (anonymous):

thanks :) @UsukiDoll